# Supershear–subshear–supershear rupture sequence during the 2025 Mandalay Earthquake in Myanmar

### v2

@@ 3 Discussion
To investigate the reason for the abrupt deceleration of the rupture, we calculated the surface fault slip distribution (Fig.2b) using satellite imagery. North-south surface displacements obtained at 80-m resolution reveal that slip occurred precisely along the known fault trace (Figure S1). By taking their difference between the east and west, we can derive the surface slip distribution. The results show that the slip reached 5 m near the epicenter. However, it exhibits a local minimum of 2–3 m at $\sim 50$ km south of the epicenter. Further south, a widespread slip of 4 m is distributed over an epicentral distance of about 100–250 km (see Supplementary Material for details).
@@
Considering that fault slip distribution is generally proportional to the static stress drop, it can be inferred that a large static stress drop occurred near the epicenter. Conversely, there likely existed a region with a small stress drop, centered $\sim 50$ km from the epicenter ($\sim 21.5^{\circ}$). Therefore, while the rupture accelerated to the P-wave speed near the epicenter due to the large stress drop, the energy release gradually weakened as the rupture propagated around the 50-kilometer point, making it unable to maintain supershear. Theoretically, rupture propagation between the Rayleigh wave speed and $\sqrt{2}$ times the S-wave speed is unstable because the singularity is weaker than the square root; consequently, the energy release rate decreases as the rupture decelerates within this speed range [6]. Hence, the rupture might immediately revert to subshear regime due to the positive feedback between the deceleration of the propagation speed and reduction of energy release rate. Subsequently, beyond the 100-kilometer point, an increase in both stress drop and energy release may have led to the re-establishment of supershear rupture.
@@
The local minimum of the released potency density was also independently identified by a teleseismic inversion analysis [10]. Some structures that may behave as a barrier and reduce the slip at $\sim 21.5^{\circ}$ on the Sagaing fault have been previously suggested. Some investigations have revealed that the location $\sim 21.5^{\circ}$ is a fault segment boundary characterized by the edge of a seismic gap and a discontinuity in the fault dip angle [9, 20]. Our model may thus provide a case study of the interaction between a dramatic change in rupture speed and a geological barrier.
@@ 4 Conclusions
This study provides unprecedented insights into the rupture characteristics of the 2025 Mandalay Earthquake (Mw 7.7) along the Sagaing Fault. Through the integrated analysis of a unique coseismic video recording, seismic data, and satellite imagery, we have revealed a dynamic rupture history.
Our findings indicate that the rupture initiated with supershear velocities near the hypocenter but experienced a notable deceleration to subshear speeds as it approached the CCTV observation site. This temporary slowdown is strongly supported by observed fault-normal acceleration patterns and aligns with the presence of a localized minimum in fault slip around 50 km south of the epicenter, suggesting a zone of reduced stress drop. Subsequently, beyond this point, the rupture appears to have re-established its supershear propagation.
@@ acknowledgements
This study was supported by ERI JURP 2025-S-B102. This study was conducted using the FUJITSU Supercomputer PRIMEHPC FX1000 and FUJITSU Server PRIMERGY GX2570 (Wisteria/BDEC-01) at the Information Technology Center, the University of Tokyo. The authors thank Dr. Ryo Okuwaki for their helpful comments.
@@
The source code for 2-D BEM calculation is available at https://doi.org/10.17605/OSF.IO/MTR76. OpenSWPC, the finite difference code, for our 3-D calculation is available at https://doi.org/10.5281/zenodo.15665702. The surface displacement in Supplementary Material is based on Copernicus Sentinel data (2025) processed by Sentinel Hub (https://dataspace.copernicus.eu/explore-data/data-collections/sentinel-data/sentinel-2).
@@ Appendix A Supplementary Material: Methods
A.1 2-D kinematic rupture simulation by a boundary element method
We calculated on-fault acceleration during the rupture propagation by 2-D boundary element method. The numerical integration is nothing but calculating equation (32) of Tada & Madariaga[19]. In their notation, the flat 2-D crack is along the $x_{1}$-axis. We consider strike-slip faulting, where the slip direction and fault-normal direction are along the $x_{1}$- and $x_{2}$-directions, respectively. Hence, our equation is
$\dot{u}_{2}^{i,j,n}=\sum_{k,m}D_{1}^{k,m}\,\dot{K}_{1,u2}^{i,j,k,n,m},$ (S.1)
and we obtain the acceleration waveform by differentiating eq.(S.1) with respect to time.
@@
where $H(\cdot)$ is the Heaviside function, and $t_{m}$ is the $m$-th time collocation point. $t_{r}^{k}$ is the rupture arrival time at $x=x_{k}$, where $x_{k}$ is the $k$-th discretized fault segment edge. For Figure 3a and 3b in the main text, we employed the following two scenarios, respectively:
  • If the rupture velocity is 5.125 km/s (constant) everywhere, then $t_{r}^{k}=\dfrac{x_{k}}{5.125}$.
  • If rupture velocity is 3 km/s in 82 km $<x<$ 124 km and is 6 km/s otherwise, then $t_{r}^{k}=\begin{cases}\frac{1}{6}x_{k}&(0\leq x_{k}\leq 82)\\ \frac{82}{6}+\frac{1}{3}(x_{k}-82)&(82<x_{k}<124)\\ \frac{82}{6}+\frac{1}{3}(124-82)+\frac{1}{6}(x_{k}-124)&(124\leq x_{k})\\ \end{cases}$.
In the second scenario, the points where the rupture velocity jumps and the subshear rupture velocity (3 km/s) are chosen such that $t_{r}^{k}=48$ s at $x_{k}=246$ km, satisfying the rupture arrival condition at NPW station.
A.2 3-D kinematic rupture simulation by a finite difference method
@@
The equations of motion and the constitutive relations based on the generalized Zener body were numerically solved using the three-dimensional parallel finite difference code OpenSWPC[14, 13]. A three-dimensional viscoelastic medium with dimensions of 500 km (north-south) $\times$ 200 km (east-west) $\times$ 100 km (vertical) was discretized using a uniform grid with 100-meter spacing in each direction. The medium was assumed to be a homogeneous, half-infinite viscoelastic body with a mass density of 2.7 g/cm3, P-wave velocity of 6.0 km/s, S-wave velocity of 3.5 km/s, and intrinsic attenuation factors of Qp = 600 and Qs = 300. The ground surface was defined at 5 km from the top of the domain, and traction-free boundary conditions[16] were applied there. The earthquake source was represented as a collection of discrete point sources. A rectangular region in the central part of the model, extending 300 km in the north-south direction and 20 km in depth from the ground surface, was designated as the fault plane. Stress changes associated with fault slip[4, 17] were applied to the numerical grid points corresponding to the fault for the duration of the rise time starting from the rupture onset time. The rupture onset time at each point was computed according to the setup described in the main text, with rupture initiation assumed at the northern end of the fault. To suppress artificial reflections of seismic waves from the model boundaries, the ADE-CFS perfectly matched layer (PML) scheme[22] with a thickness of 20 grid points was applied along the outer edges of the computational domain. The model was partitioned into 32 $\times$ 32 segments in the horizontal directions, and the simulation was executed on 256 CPUs. Four MPI processes were launched per CPU, with each process utilizing one-fourth of the CPU’s cores via OpenMP threading. The computation of 60,000 time steps with a time increment of 0.005 s took approximately 30 minutes.
A.3 Estimation of Surface fault slip
@@
To estimate the coseismic surface fault slip, we performed pixel-offset (optical correlation) analysis using Sentinel-2 satellite imagery obtained before and after the earthquake. Here, we used 10 m-resolved Band-3 images of the L2A product, with a central wavelength of 560 nm, obtained on March 20th and 30th, 2025.
@@
For the pixel-offset analysis, we used geoCosiCorr3D [1, 2]. Here, we employed 32 pixels (320 m) as a window size for correlating the images, and estimated northward and eastward displacement for every eight steps (80 m). Estimated surface displacements were mapped in Figure S.4.
@@
The distribution of surface fault slip (Figure 2b in the main text) was estimated from the discrepancies in northward components between both sides of the surface fault trace, assuming that this component corresponds with the fault displacements, as the fault strike is almost north-south, and the east-west component has hardly been detected. Considering the window size of the pixel-offset analysis, we excluded 1 pixel on the fault trace and 2 pixels (160 m) on each side from the fault in the analysis results. We then calculated the average displacements ranging from 3 to 12 pixels (800 m) on each side of the fault on the east-west cross-sections.
@@
Figure S.4: Surface displacement estimated by the pixel-offset (optical correlation) analysis of Sentinel-2 imagery. (a) Northward and (b) Eastward displacements. The red star and the X mark show the locations of the epicenter and CCTV, respectively. Fault traces are based on [18].

### v3

@@ 3 Discussion
To investigate the reason for the abrupt deceleration of the rupture, we calculated the surface fault slip distribution (Fig.2b) using satellite imagery. North-south surface displacements obtained at 80-m resolution reveal that slip occurred precisely along the known fault trace (Fig.S3a). By taking their difference between the east and west, we can derive the surface slip distribution (see Supplementary Material for details). The results show that the slip reached 5 m near the epicenter. However, it exhibits a local minimum of 2–3 m at 40–60 km south of the epicenter. Further south, a widespread average slip of 4 m, without a broad region of reduction, is distributed over an epicentral distance of approximately 100–250 km.
Final slip amount correlates with the static stress drop. Mathematically, the slip distribution $D(\boldsymbol{k})$ and stress drop distribution $\tau(\boldsymbol{k})$ on a planar fault are related as $D(\boldsymbol{k})\sim-\dfrac{2}{\mu|\boldsymbol{k}|}\tau(\boldsymbol{k})$, where $\boldsymbol{k}$ is the 2-D wavenumber vector (Andrews,, 1980). This indicates that $D$ is obtained by reducing the shorter wavelength (i.e., larger $|\boldsymbol{k}|$) components of $\tau$. This relationship, resulting from numerical simulations with heterogeneous stress drop, is also evident in Figs. 3 and 4 of Andrews & Barall, (2011). Considering this, a large static stress drop can be inferred near the epicenter of the Myanmar earthquake. Conversely, a region with a small stress drop likely existed, centered 40–60 km from the epicenter ($\sim 21.5^{\circ}$).
@@
In general, local rupture acceleration stems from the energy release rate, which scales with the square of the stress intensity factor. The stress intensity factor itself scales with stress drop (Freund 1990). This indicates a theoretical local correlation between stress drop and rupture velocity. Therefore, although the rupture accelerated to supershear speed near the epicenter due to the large stress drop, energy release gradually weakened around the 40–60 kilometer point, preventing it from maintaining supershear propagation. Theoretically, rupture propagation between the Rayleigh wave speed and $\sqrt{2}$ times the S-wave speed is unstable because the stress singularity is weaker than the square root (Freund,, 1990, section 4.3.4). Consequently, the energy release rate decreases as the rupture decelerates within this speed range. Hence, the rupture might immediately revert to subshear regime due to the positive feedback between rupture deceleration and energy release rate reduction. Subsequently, beyond the 100-kilometer point, increased stress drop and energy release may have led to the re-establishment of supershear rupture.
@@
For the location $\sim 21.5^{\circ}$, Inoue et al., (2025) revealed the local minimum of the released potency density and the variation of dipping angle of the fault by their potency density tensor inversion method that enables to image not only slip amount but also strike/dip/rake angles on each point along the fault plane. Also, some investigations have revealed that the location $\sim 21.5^{\circ}$ is a fault segment boundary characterized by the edge of a seismic gap and a discontinuity in the fault dip angle (Hurukawa & Maung Maung,, 2011; Tha Zin Htet Tin et al.,, 2022). These seismological and geological results suggest that there exists a segment boundary that may behave as a barrier and reduce the slip at $\sim 21.5^{\circ}$. Our model may thus provide a case study of the interaction between a dramatic change in rupture speed and a geological barrier.
@@ 4 Conclusions
This study provides unprecedented insights into the rupture characteristics of the 2025 Mandalay Earthquake ($M_{w}$7.7) along the Sagaing Fault. Through the integrated analysis of a unique coseismic video recording, seismic data, kinematic simulation, and satellite imagery, we have revealed a dynamic rupture history.
Our findings indicate that the rupture initiated with supershear velocities near the hypocenter but experienced a notable deceleration to subshear speeds as it approached the CCTV observation site. This temporary slowdown is strongly supported by observed fault-normal acceleration patterns and aligns with the presence of a localized minimum in fault slip around 40–60 km south of the epicenter, suggesting a zone of reduced stress drop. Subsequently, beyond this point, the rupture appears to have re-established its supershear propagation.
@@ acknowledgements
This study was supported by ERI JURP 2025-S-B102. This study was conducted using the FUJITSU Supercomputer PRIMEHPC FX1000 and FUJITSU Server PRIMERGY GX2570 (Wisteria/BDEC-01) at the Information Technology Center, the University of Tokyo. Thorough reviews by Dr. Kiran Kumar Thingbaijam and Dr. Tiegan Hobbs contributed to improving the manuscript. The authors thank Dr. Ryo Okuwaki for their helpful comments.
@@
The video, credited to GP Energy Myanmar, was originally uploaded to Facebook (https://www.facebook.com/htin.aung.33/videos/1041579804084512) on May 11, 2025, and later reposted on YouTube (https://www.youtube.com/watch?v=77ubC4bcgRM), where the CCTV site’s coordinates were first disclosed. The significant portion of the video is featured in CBC video news (https://www.cbc.ca/player/play/video/9.6763872). The authors believe that using the video in this manuscript is consistent with the principles of fair use under U.S. copyright law for the purposes of scholarship and research (https://www.copyright.gov/fair-use/). All links in this paragraph were last accessed on July 15, 2025.
The source code for 2-D BEM calculation is available at https://doi.org/10.17605/OSF.IO/MTR76. OpenSWPC, the finite difference code, for our 3-D calculation is available at https://doi.org/10.5281/zenodo.16264099. The surface displacement in Supplementary Material is based on Copernicus Sentinel data (2025) processed by Sentinel Hub (https://dataspace.copernicus.eu/explore-data/data-collections/sentinel-data/sentinel-2).
@@ Appendix A Supplementary Material: Methods
A.1 Image analysis to extract slip history from the video
In the initial analysis of the video (Hirano,, 2025), the author executed the following steps as shown in Fig.S.1:
  • Step 1: A 100$\times$20 px region, starting 152 px from the top and 930 px from the left, in video frames was cropped from each video frame.
  • Step 2: The cropped images were compressed vertically to 100$\times$1 px, to obtain a vertical average.
  • Step 3: Edges in the 1-px-height images were enhanced.
  • Step 4: The distance between the leftmost and rightmost white pixels was measured.
A.2 Calculations to obtain the deceleration point and subshear rupture velocity
The initial supershear rupture velocity $V_{\text{sup}}$ falls between $V_{P}$ (= 6.0 km/s) and $V_{S}$ (= 3.5 km/s). At the CCTV site, from the video frames, the P-wave duration $\tau_{P}$ is, though uncertain, at least 3 seconds, while S-wave duration $\tau_{S}$ is approximately 2 seconds. The S-wave front arrival time therein is $t_{C}^{S}:=V_{P}^{-1}r_{C}+\tau_{P}$, and the equation of the front (gray dotted line in Fig.S.2a) is given by
$r=V_{S}\left(t-t_{C}^{S}\right)+r_{C}.$
The intersection of this line and the initial rupture propagation represented by $r=V_{\text{sup}}t$ gives the deceleration point which is the origin of the observed S-wave. The solution $r_{d}$ is calculated as:
$r_{d}=$ $\frac{r_{C}-V_{S}t_{C}^{S}}{1-V_{S}/V_{\text{sup}}}.$
The subshear rupture velocity $V_{\text{sub}}$ from the deceleration point to CCTV is
$V_{\text{sub}}=$ $\frac{r_{C}-r_{d}}{t_{C}^{S}+\tau_{S}-r_{d}/V_{\text{sup}}},$
where $t_{C}^{S}+\tau_{S}$ and $r_{d}/V_{\text{sup}}$ in the denominator represent the rupture times at the CCTV and the deceleration point, respectively. From the above, unknown parameters are $\tau_{P}$ and $V_{\text{sup}}$. Given that rupture reached $r=r_{N}$ at $t=t_{N}$, the following condition must be satisfied:
$t_{C}^{S}+\tau_{S}\leq$ $t_{N}-\frac{r_{N}-r_{C}}{V_{P}}.$ (S.1)
where the right-hand side denotes the boundary between the white and top-right gray regions in Fig.S.2a. Otherwise, the rupture cannot reach NPW at $t=t_{N}$, even if it propagates at $V_{P}$ after the CCTV site.
We can read $t_{N}=48$ from the Fig.4 of Lai et al., (2025), who assumed that the origin time was 2025-03-28 06:20:54 UTC, though a rationale was not provided. This origin time is 2 seconds behind the USGS origin time, suggesting that $t_{N}=50$ s may be more appropriate if the USGS time is adopted. If $V_{\text{sup}}=V_{P}=6$ km/s, the deceleration point and subshear rupture velocity for the earliest ($\tau_{P}=3$ s) and latest (the equality in (S.1) holds) cases are $(r_{d},V_{\text{sub}})=(98.8,2.74)$ and =(65.2,3.13)=(65.2,3.13)= ( 65.2 , 3.13 ), respectively. Fig.S.2b indicates that $V_{\text{sup}}\leq 5$ is difficult to achieve, as $V_{\text{sub}}$ becomes excessively slow, particularly in the earliest case. Furthermore, $V_{\text{sup}}<\sqrt{2}V_{S}\sim 4.95$ is unstable, as explained in the Discussion section. For numerical simulations, the latest case was employed because $r_{d}=65.2$ is more comparable to the local minimum of the surface slip as in Discussion section.
A.3 2-D kinematic rupture simulation by a boundary element method
We calculated on-fault acceleration during the rupture propagation by 2-D boundary element method. The numerical integration is performed by calculating equation (32) of Tada & Madariaga, (2000). In their notation, the flat 2-D crack is along the $x_{1}$-axis. We consider strike-slip faulting, where the slip direction and fault-normal direction are along the $x_{1}$- and $x_{2}$-directions, respectively. Hence, our equation is
$\dot{u}_{2}^{i,j,n}=\sum_{k,m}D_{1}^{k,m}\,\dot{K}_{1,u2}^{i,j,k,n,m},.$ (S.2)
Differentiating eq.(S.2) with respect to time yields the acceleration waveform.
@@ A.2 Calculations to obtain the deceleration point and subshear rupture velocity
where $H(\cdot)$ is the Heaviside function, and $t_{m}$ is the $m$-th time collocation point. $t_{r}^{k}$ is the rupture arrival time at $x=x_{k}$, where $x_{k}$ is the $k$-th discretized fault segment edge. For Fig.3a and 3b in the main text, we employed the following two scenarios, respectively, with some parameters obtained in A.2:
  • If the rupture velocity is 4.92 km/s (constant) everywhere, then $t_{r}^{k}=\dfrac{x_{k}}{4.92}$.
  • If rupture velocity is 3.13 km/s in 65.2 km $<x<$ 124 km and is 6 km/s otherwise, then $t_{r}^{k}=\begin{cases}\frac{1}{6}x_{k}&(0\leq x_{k}\leq 65.2)\\ \frac{65.2}{6}+\frac{1}{3}(x_{k}-65.2)&(65.2<x_{k}<124)\\ \frac{65.2}{6}+\frac{1}{3}(124-65.2)+\frac{1}{6}(x_{k}-124)&(124\leq x_{k})\\ \end{cases}$.
A.4 3-D kinematic rupture simulation by a finite difference method
@@
We numerically solved the equations of motion and constitutive relations for a generalized Zener body using the 3-D parallel finite difference code OpenSWPC (Maeda et al.,, 2017; Maeda,, 2025). A 3-D medium with dimensions of 500 km (north-south) $\times$ 200 km (east-west) $\times$ 100 km (vertical) was discretized using a uniform grid with 100-meter spacing in each direction. The medium was assumed to be a homogeneous, viscoelastic half-space with a mass density of 2.7 g/cm3, P-wave velocity of 6.0 km/s, S-wave velocity of 3.5 km/s, and intrinsic attenuation factors of Qp = 600 and Qs = 300. The ground surface was defined at 5 km from the top of the domain, and traction-free boundary conditions (Nakamura et al.,, 2012) were applied there. The earthquake source was represented as a collection of discrete point sources. A rectangular region in the central part of the model, extending 300 km in the north-south direction and 20 km in depth from the ground surface, was designated as the fault plane. Stress changes associated with fault slip (Coutant et al.,, 1995; Pitarka,, 1999) were applied to the numerical grid points corresponding to the fault for the duration of the rise time starting from the rupture onset time. The rupture onset time at each point was computed according to the setup described in the main text, with rupture initiation assumed at the northern end of the fault. To suppress artificial reflections from the model boundaries, the ADE-CFS perfectly matched layer scheme (Zhang & Shen,, 2010) with a thickness of 20 grid points was applied along the outer edges of the computational domain. The model was partitioned into 32 $\times$ 32 segments in the horizontal directions, and the simulation was executed on 256 CPUs. Four MPI processes were launched per CPU, with each process utilizing one-fourth of the CPU’s cores via OpenMP threading. The computation of 60,000 time steps with a time increment of 0.005 s took approximately 30 minutes.
A.5 Estimation of Surface fault slip
@@
We estimated coseismic surface fault slip by pixel-offset (optical correlation) analysis of pre- and post-seismic Sentinel-2 satellite imagery. Here, we used 10 m-resolved Band-3 images of the L2A product, with a central wavelength of 560 nm, obtained on March 20th and 30th, 2025.
@@
For the pixel-offset analysis, we used geoCosiCorr3D (Aati et al., 2022a, ; Aati et al., 2022b, ). Here, we employed 32 pixels (320 m) as a window size for correlating the images, and estimated northward and eastward displacement for every eight steps (80 m). Estimated surface displacements were mapped in Fig.S.3.
@@
The distribution of surface fault slip (Fig.2b in the main text) was estimated from the discrepancies in northward components between both sides of the surface fault trace, assuming that this component corresponds to the fault displacements, as the fault strike is almost north-south, and the east-west component was negligible. Considering the window size of the pixel-offset analysis, we excluded 1 pixel on the fault trace and 2 pixels (160 m) on each side from the fault in the analysis results. We then calculated the average displacements ranging from 3 to 12 pixels (800 m) on each side of the fault on the east-west cross-sections.
Figure S.1: Steps for image analysis to quantify slip history on the surface fault. The leftmost white pixel represents a part of the gate structure in front of the fault, while the rightmost one, a part of a pole on the opposite side.
Figure S.2: (a) Schematic illustration of possible rupture front to CCTV $(r=r_{C})$. The rupture front, with a velocity of up to $V_{P}$ from the origin to the black triangle (NPW), is unable to penetrate the gray-filled regions. At CCTV, the gray $(\tau_{P})$ and the black $(\tau_{S})$ horizontal bars correspond to the durations of the P-wave and S-wave, respectively. The deceleration point of the rupture front, $r_{d}$, is the intersection of the early supershear rupture front and the S-wave. (b) Subshear rupture velocity as a function of the supershear rupture velocity, $V_{\text{sup}}$. The earliest case corresponds to $\tau_{P}=3$ s, while the end of $\tau_{S}$ touches the forbidden region in the latest case. The vertical dashed line in (b) represents $\sqrt{2}V_{S}$ (see the main text for its meaning).
@@
Figure S.3: Surface displacement estimated by the pixel-offset (optical correlation) analysis of Sentinel-2 imagery. (a) Northward and (b) Eastward displacements. The red star and the $\times$ mark show the locations of the epicenter and CCTV, respectively. Fault traces are based on Styron & Pagani, (2020).